A race car accelerates uniformly from 18.5 m/s to 23.6m/s in 21seconds. Determine the acceleration of the car and the distance traveled





I need this asap

1
Expert's answer
2022-05-16T13:40:53-0400
"a=\frac{\Delta v}{\Delta t}=\frac{23.6-18.5}{21}=0.24\:\rm m/s^2"